Determine the amount of volume of 37 wt% HCl that should be added to 1 L of distilled water to make 0.1 M HCl and prepare the solution. The density should be obtained for HCl to determine the volume needed.
Answer #1

Solution:

The molar mass of HCl =. 36.4611 g/mol
Density of HCl = 1.189 g/ml

According to Molarity equation,

Molarity (M) =mass of HCl / molar mass of HCl x volume in L

Mass of HCl = 0.1 M x 36.4611 g/mol x 1 Litre = 3.6461 g

Weight = 37 % = 37 /100 = 0.370

Volume of HCl needed = mass of HCl / density x weight of HCl

Volume of HCl = (3.6461g) /(0.370 x 1.189g /ml) = 8.2879 ml

So, 8.2879 ml of 37% concentrated HCl is needed to prepare 0.1 M HCL.